# Human-in-the-Loop Causal Discovery under Latent Confounding using Ancestral GFlowNets

### v2

@@ 3 Ancestral GFlowNets
We propose AGFN, a GFlowNet-based method for sampling AGs using a score function. Specifically, AGFN encompasses a GFlowNet with the following characteristics:
@@
[…]
Action constraints. To ensure AGFN only samples AGs, we mask out actions that would lead to paths forming cycles or almost cycles. To achieve this, we verify whether the resulting graph respects Bhattacharya et al. (2021)’s algebraic characterization of the space of AGs. More specifically, any AG $\mathcal{G}$ is characterized by an adjacency matrix $\mathbf{A}_{d}\in\mathbb{R}^{n\times n}$ for its directed edges and another adjacency matrix $\mathbf{A}_{b}\in\mathbb{R}^{n\times n}$ for its bidirected edges, adhering to:
$\text{trace}(\exp\{\mathbf{A}_{d}\})-n+\mathbf{1}^{T}(\exp\{\mathbf{A}_{d}\}\odot\mathbf{A}_{b})\mathbf{1}=0,$ (8)
in which $\mathbf{1}$ is a $n$-dimensional unit vector and $\odot$ denotes the Hadamard (elementwise) product of matrices.
Score-based belief. We propose using a strictly decreasing transformation of a score function $U$ as the reward $R$ for AGFN. More precisely, we define $R$ as
$R(\mathcal{G})=\exp\left\{\frac{\mu-U(\mathcal{G})}{\sigma}\right\}$ (9)
for given constants $\mu\in\mathbb{R}$ and $\sigma\in\mathbb{R}^{+}$ that ensure numerical stability (Zhang et al., 2023). In practice, we sample $S$ AGs $\{\mathcal{G}^{(s)}\}_{s=1}^{S}$ from an untrained AGFN, and set $\mu=\nicefrac{{1}}{{S}}\sum_{s}U(\mathcal{G}^{(s)})$ and $\sigma=\sqrt{\nicefrac{{1}}{{S}}\sum_{s}(U(\mathcal{G}^{(s)})-\mu)^{2}}$.
Score for linear Gaussian models. Since we focus on linear Gaussian models, we choose the extended Bayesian Information Criterion Foygel and Drton (2010) as our score function. Specifically, for any AG $\mathcal{G}=(\mathbf{V},\mathbf{E})$:
in which $(\hat{\mathbf{B}},\hat{\boldsymbol{\Omega}})$ is the MLE estimate of model parameters (see eq. 3) obtained using the residual iterative conditional fitting algorithm Drton et al. (2009).
Forward flow. We use a Graph Isomorphism Network (GIN) Xu et al. (2019) $\Phi$ to compute a $d$-dimensional representation for each node in the AG $\mathcal{G}_{t}$ at the $t$-th step of the generative process and use sum pooling to get an embedding for $\mathcal{G}_{t}$. Then, considering $\mathcal{A}_{t}$ as the space of feasible actions at $\mathcal{G}_{t}$ (i.e., those leading to an AG), we use an MLP $\phi\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\mathcal{A}_{t}|}$ with a softmax activation at its last layer to map $\mathcal{G}_{t}$’s embedding to a distribution over $\mathcal{A}_{t}$. More precisely, given $\mathbf{H}^{(t)}=\Phi(\mathcal{G}_{t})\in\mathbb{R}^{|\mathbf{V}|\times d}$, we compute
$\mathbf{p}=\phi\left(\sum_{v\in\mathbf{V}}\mathbf{H}_{v}^{(t)}\right)$ (11)
as the probability distribution over the feasible actions at $\mathcal{G}_{t}$.
Backward flow. Backward actions correspond to removing edges. Following Shen et al. (2023), we parametrize the backward flow $F_{B}$ with an MLP and alternate between updating $\pi_{F}$ and $\pi_{F_{B}}$, using gradient-based optimization.
4 Human-in-the-Loop Causal Discovery
Concluding the training, we propose leveraging the AGFN-generated samples to design questions for the expert that optimize the reduction of entropy in the distribution $p_{\theta}(\mathcal{G})$ over the space of AGs. Then, we use the human feedback to update $p_{\theta}(\mathcal{G})$, and iteratively repeat the process. The following paragraphs describe i) how we model human feedback, ii) how we update our belief over AGs given human responses, and iii) our experimental design strategy for expert inquiry.
Modeling human feedback.
We assume humans are capable of answering questions regarding the ancestral relationship between pairs of random variables. In this case, we model their prior knowledge on a relation $r=\{U,V\}$ between nodes $U,V\in\mathbf{V}$ as a categorical distribution over a random variable denoted $\omega_{r}$. Fix an arbitrary total order $<$ in $\mathbf{V}$. By definition, $\omega_{r}=1$ if there is no edge between $U$ and $V$; $\omega_{r}=2$ if $U$ is ancestor of $V<U$; $\omega_{r}=3$ if $V$ is ancestor of $U>V$; and $\omega_{r}=4$ if there is a bidirected edge between $U$ and $V$. Since the human has access to our AGFN before being probed for the first time, we set $\rho_{r,k}=p_{\theta}(\omega_{r}=k)$ as the prior probability of $\omega_{r}=k$. Moreover, we consider that the expert’s feedback $f_{r}\in\{1,2,3,4\}$ on the relation $r$ is a noisy realization of the true, unobserved value of the relation feature $\omega_{r}$ under the expert’s model. Putting all elements together results in a two-level Bayesian hierarchical scheme for categorical data:
@@ 4 Human-in-the-Loop Causal Discovery
with $\eta_{r}=\rho_{r,f_{r}}\cdot\pi+\left(\frac{1-\pi}{3}\right)\cdot(1-\rho_{r,f_{r}})$.
Figure 3: Sampling quality. The reward-induced marginal distribution over graph features is adequately approximated by the marginal distribution learned by the GFlowNet.
Updating beliefs.
We update our AGFN by weighing it by our posterior over the expert’s knowledge, described in the previous paragraph, similarly to a product-of-experts approach Hinton (2002). For this, let $\mathbf{f}_{K}=(f_{r_{k}})_{1\leq k\leq K}$ be the sequence of $K$ feedbacks issued by the expert and define our novel belief distribution $q(\mathcal{G};\mathbf{f}_{K})$ over the space of AGs as
$q(\mathcal{G};\mathbf{f}_{K})\propto p_{\theta}(\mathcal{G})\prod_{1\leq k\leq K}p(\omega_{r_{k}}|f_{r_{k}}).$ (15)
Importantly, we use $p_{\theta}$ as proposal distribution in importance (re-)sampling Gordon et al. (1993) to approximately sample from $q(\mathcal{G};\mathbf{f}_{K})$ — or to approximate the expected value of a test function. More precisely, we estimate the value of a function $h$ over the space of AGs as:
$\mathbb{E}_{q}[h(\mathcal{G})]\approx\sum_{t=1}^{T}c^{-1}\frac{q(\mathcal{G}^{(t)};\mathbf{f}_{K})}{p_{\theta}(\mathcal{G}^{(t)})}h(\mathcal{G}^{(t)})$ (16)
with $(\mathcal{G}^{(t)})_{t=1}^{T}\sim p_{\theta}(\mathcal{G})$ and $c=\sum_{t=1}^{T}\nicefrac{{q(\mathcal{G}^{(t)};\mathbf{f}_{K})}}{{p_{\theta}(\mathcal{G}^{(t)})}}$.
Active knowledge elicitation.
@@
To make the most out of possibly costly human interactions, we query the human about the relation that maximally reduces the expected cross-entropy between our belief over AGs before and after human feedback. More precisely, we define an acquisition function $a_{k}:{{\mathbf{V}}\choose{2}}\rightarrow\mathbb{R}$ for the $k>1$-th inquiry as:
$a_{k}(r)=-\mathbb{E}_{f_{r}\sim p(\cdot|\mathbf{f}_{K})}\big{[}\mathbf{H}\left(q(\mathcal{G};\mathbf{f}_{K},f_{r}),q(\mathcal{G};\mathbf{f}_{K})\right)\!\big{]}$ (17)
@@
in which $p(f_{r}|\mathbf{f}_{K})$ is the posterior predictive distribution according to the user model, $q_{0}\propto R$ and $\mathbf{H}(\cdot,\cdot)$ is the cross-entropy. Then, we maximize this acquisition to select which relation $\tilde{r}_{k}$ we will probe the expert, i.e.:
$\tilde{r}_{k}=\operatorname*{arg\ max}_{r\in{{\mathbf{V}}\choose{2}}}a_{k}(r).$ (18)
@@
As aforementioned, we use importance sampling with $q_{0}$ as a proposal to estimate the acquisition function $a_{k}$. This allows us to leverage AGFN samples and effectively avoid the need for retraining them. It is worth mentioning that because $\mathbf{H}(p,p^{\prime})\geq\mathbf{H}(p,p)$ for any two distributions $p$ and $p^{\prime}$ of the same support, our strategy is equivalent to minimizing an upper bound on the entropy of $q_{k}$. Also, different from acquisitions based on information gain or mutual information, approximating Equation 17 via Monte Carlo does not require exhaustive integration over the space of AGs to yield asymptotically unbiased estimates — see Appendix.
@@
SHD BIC SHD BIC SHD BIC
FCI $3.03{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.13}}$ 1.13 $5481.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.69}}$ 2.69 $\mathbf{3.75}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.64}}$ 0.64 $\mathbf{5426.18}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.74}}$ 1.74 $6.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.20}}$ 1.20 $5433.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 6.94}}$ 6.94
GFCI $\mathbf{2.24}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.64}}$ 0.64 $\mathbf{5479.77}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.75}}$ 1.75 $4.21{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.96}}$ 0.96 $5427.09{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.85}}$ 2.85 $\mathbf{5.23}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.08}}$ 1.08 $\mathbf{5431.67}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 7.91}}$ 7.91
DCD $3.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.30}}$ 1.30 $5482.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 5.16}}$ 5.16 $5.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.23}}$ 1.23 $5429.51{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 4.37}}$ 4.37 $6.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.22}}$ 1.22 $5436.84{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 9.41}}$ 9.41
@@
AGFN (ours) $\mathbf{6.04}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.12}}$ 2.12 $\mathbf{5494.67}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 37.08}}$ 37.08 $8.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.04}}$ 2.04 $\mathbf{5456.16}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 52.25}}$ 52.25 $\mathbf{6.58}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.34}}$ 2.34 $\mathbf{5478.01}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 40.36}}$ 40.36
Table 1: Average SHD and BIC. The denotes methods yielding point estimates. We use Bootstrap to report the mean and average standard deviation for these. For N-ADMG and AGFN, we estimate the quantities using $100$k samples.
5 Experiments
Our experiments have three objectives. First, we validate that AGFN can accurately learn the target distribution over the space of AGs. Second, we show that AGFN performs competitively with alternative methods on three data sets. Third, we attest that our experimental design for incorporating the expert’s feedback efficiently reduces the uncertainty over AGFN’s distribution. We provide further experimental details in the Appendix. Code is available in the supplement.
5.1 Distributional Assessment of AGFN
Data. Since violations of faithfulness are more likely in dense graphs Uhler et al. (2012), we create $20$ $5$-node random graphs Uhler et al. (2012) from a directed configuration model (Newman, 2010) whose in- and out-degrees are uniformly sampled from $\{0,1,2,3,4\}$. We draw $500$ independent samples from a structure-compatible linear Gaussian SCM with random parameters for each graph.

### v3

@@ 3 ANCESTRAL GFLOWNETS
To quantify uncertainty in CD under latent confounding and enhance inferences via active learning, we propose AGFN, a GFlowNet-based probabilistic method that samples AGs from a score-based belief distribution. AGFN is built around a GFlowNet in which:
@@
The generation of a trajectory $\tau=\{\{\},\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{T}\}$ begins with an empty graph with nodes $\mathbf{V}$, iteratively adding edges of types $\{\leftarrow,\rightarrow,\leftrightarrow\}$ between pairs of variables. The following sections describe AGFN. For additional details refer to the Appendix.
3.1 Action constraints
To ensure AGFN samples only valid AGs, we restrict actions that would create cycles or almost cycles. An action adding a directed edge leads to a cycle (resp. almost cycle) if and only if a directed path (resp. directed path with a single bidirected edge) exists between its endpoints, with a similar reasoning for adding bidirected edges. Similarly to Giudici and Castelo, (2003) and (Deleu et al.,, 2022, Appendix C), we leverage the iterative nature of the generative process to efficiently track such paths and mask invalid actions, as discussed in Appendix A.
3.2 Score-based belief
We define AGFN’s reward $R$ as a strictly increasing transformation of a score function $U$ measuring the goodness-of-fit of an AG to the observed data. To ensure numerical stability (Zhang et al.,, 2023), we choose constants $\mu\in\mathbb{R}$ and $\sigma\in\mathbb{R}^{+}$ and define the reward function $R(\cal{G})$ for the AG $\cal{G}$ as
$R(\mathcal{G})=\exp\left\{\frac{U(\mathcal{G})-\mu}{\sigma}\right\}.$ (8)
In practice, we sample $S$ AGs, $\{\mathcal{G}^{(s)}\}_{s=1}^{S}$, from an untrained AGFN, and define
$\mu=\frac{1}{S}\sum_{s}U(\mathcal{G}^{(s)});\,\,\sigma=\sqrt{\frac{1}{S}\sum_{s}(U(\mathcal{G}^{(s)})-\mu)^{2}}.$ (9)
Drawing on the generative process described in Section 3.1 and the reward function defined above, we learn an AGFN by minimizing the loss function in Equation (7) through stochastic gradient descent (Kingma and Ba,, 2017). For further details regarding AGFN training and the parametrization of the GFlowNet’s policies, please consult Section C.2.1.
4 EITL Causal Discovery
After training, we leverage AGFN-generated samples to design expert queries and refine the learned distribution based on their answers. Queries are selected to maximally reduce the entropy of the distribution $p_{\theta}(\mathcal{G})$ over AGs. Expert feedback is then used to update $p_{\theta}(\mathcal{G})$, and the process is repeated. Next we describe i) how feedback is modeled, ii) how beliefs over AGs are updated given expert responses, and iii) our experimental design strategy for expert queries.
4.1 Modeling expert feedback
We leverage expert feedback to determine the existence of ancestral relationships between nodes. A detailed discussion of this type of knowledge is in Section 4.5. We model prior knowledge on a relation $r=\{V_{i},V_{j}\}$ between nodes $V_{i},V_{j}\in\mathbf{V}$ as a categorical distribution over a random variable denoted $\omega_{r}$. For notational convenience, we let $\omega_{r}=1$ if there is no edge between $V_{i}$ and $V_{j}$; $\omega_{r}=2$ if $V_{i}$ is ancestor of $V_{j}<V_{i}$; $\omega_{r}=3$ if $V_{j}$ is ancestor of $V_{i}>V_{j}$; and $\omega_{r}=4$ if there is a bidirected edge between $V_{i}$ and $V_{j}$. Any other encoding for $\omega_{r}$ would yield an equivalent algorithm.
Since the expert sees AGFN before being the first query, we set $\rho_{r,k}=p_{\theta}(\omega_{r}=k)$ as the prior probability of $\omega_{r}=k$. The expert’s feedback $f_{r}\in\{1,2,3,4\}$ is treated as a noisy realization of the true, unobserved relation feature $\omega_{r}$. Together, these elements form a Bayesian hierarchical scheme for categorical data:
@@ 4.1 Modeling expert feedback
with $\eta_{r}=\rho_{r,f_{r}}\cdot\pi+\left(\frac{1-\pi}{3}\right)\cdot(1-\rho_{r,f_{r}})$ and $\odot$ representing the Hadamard product between vectors.
Figure 3: Sampling quality. AFGN accurately samples from its underlying score-based beliefs. The 1st plot (left) shows that the marginal probabilities over edges induced by AGFN match the analytical marginal. The 2nd shows the same for the probability of directed paths between two variables. The rightmost plots show that the distributions of the SHD and BIC from AGFN samples closely match the analytical one.
4.2 Updating beliefs
To understand our procedure for updating AGFN’s distribution based on the observed responses from an expert, the reader is invited to notice that the expert-induced posterior, $p(f_{r}|w_{r})$, characterizes a distribution over the edges of an AG through $q(e|\mathbf{f}_{K})\coloneqq p(w_{r}|f_{r})\cdot u(r)$. Here, $q(e)$ denotes the probability of an edge $e$ corresponding to the relation $r$ featuring $f_{r}$ (as described in Section 4.1) and $u(r)$ is an uniform distribution over the AGs’ unlabeled edges.
Similarly, an AGFN’s policy $\pi_{F}(\cdot|s)$ corresponds to a distribution over the edges of the AG $s$, which we denote by $p(e)$. From this perspective, we follow a mixture-of-experts approach (Hinton,, 2002) and let the updated AGFN’s policy be $\pi_{F}^{(K)}(s^{\prime}|s,\mathbf{f}_{K})\propto p(e)\cdot q(e|\mathbf{f}_{K})$, in which $e$ is the edge added in the transition from $s$ to $s^{\prime}$. Intuitively, this ensures that the AGFN samples graphs whose edges are consistent with both the score function and the expert’s beliefs more frequently. Additionally, the structural constraints for ancestrality of the sampled graphs remain naturally encoded into $p(e)$ (see Figure 2).
Based on the updated policy, we can easily estimate the expectation of a given test function (e.g., an acquisition function for active learning; see Section 4.3) through a Monte Carlo estimator.
4.3 Active knowledge elicitation
@@
To efficiently utilize possibly costly expert interactions, we query the relation that maximally reduces the expected cross-entropy between our belief over AGs before and after feedback. Specifically, we define an acquisition function $a_{k}:{{\mathbf{V}}\choose{2}}\rightarrow\mathbb{R}$ for the $k>1$-th inquiry as:
$a_{k}(r)=-\mathbb{E}_{f_{r}\sim p(\cdot|\mathbf{f}_{K})}\big[\mathbf{H}\left(q(\mathcal{G};\mathbf{f}_{K},f_{r}),q(\mathcal{G};\mathbf{f}_{K})\right)\!\big],$ (13)
@@
in which $p(f_{r}|\mathbf{f}_{K})$ is the posterior predictive distribution under the user model, $q_{0}\propto R$, and $\mathbf{H}(\cdot,\cdot)$ is the cross-entropy. We then maximize this acquisition to select the relation $\tilde{r}_{k}$ to query the expert, i.e.:
$\tilde{r}_{k}=\operatorname*{arg\ max}\nolimits_{r\in{{\mathbf{V}}\choose{2}}}a_{k}(r).$ (14)
@@
Since $\mathbf{H}(p,p^{\prime})\geq\mathbf{H}(p,p)$ for any distributions $p$ and $p^{\prime}$ of the same support, our strategy is equivalent to minimizing an upper bound on the entropy of $q_{k}$. Unlike acquisitions based on information gain or mutual information, Monte Carlo approximation of Eq. (13) avoids exhaustive integration over the space of AGs to yield asymptotically unbiased estimates (see Appendix C).
Algorithm 1 Simulating experts in the loop
$\{\mathcal{G}_{t}\}_{1\leq t\leq T}\text{ samples from AGFN}$, $\mathcal{G}^{*}=(\mathbf{V},E)$ true AG, $\pi$ feedback reliability
$p(\omega_{r}=k)\leftarrow\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{\omega_{r}=k\text{ in }\mathcal{G}_{t}\},\forall k\in\{1,2,3,4\},\ r\in{\mathbf{V}\choose 2}$
$\mathbf{f}\leftarrow\{\}$ $\triangleright$ Set of feedbacks (answers)
$\boldsymbol{r}\leftarrow\{\}$ $\triangleright$ Set of queries (questions)
$K\leftarrow 1$
while $\boldsymbol{r}\neq{\mathbf{V}\choose 2}$ do $\triangleright$ Iteratively query expert
  $\mathbf{r}\leftarrow\mathbf{r}\cup\{r_{K}\}$
  $\mathbf{f}\leftarrow\mathbf{f}\cup\{f_{K}\}$
  $K\leftarrow K+1$
end while
4.4 Algorithmic details
Algorithm 1 outlines our procedure for simulating expert interactions. We first estimate the marginal probabilities $p(\omega_{r}=k)$ of a relation $r$ having feature $k\in\{1,2,3,4\}$ under AGFN’s learned distribution, which serves as our prior distribution. We then iteratively select the relation that maximizes the acquisition function; the expert returns the relation’s true feature with probability $\pi$, or otherwise chooses randomly among the incorrect options.
4.5 On (non-)ancestral knowledge
Our EITL approach leverages the fact that experts can often easily judge whether one variable is an ancestor of another, without needing to determine whether the causal link is direct or confounded. Such coarse knowledge is easier to provide than the direct causation required for DAGs. For example, experimental studies frequently establish ancestrality between variables, even if they cannot rule out mediation or confounding. Likewise, experts may know that demographic or genetic variables are not downstream of diseases, or that certain associations are spurious. Our framework integrates such knowledge, along with the expert’s confidence level, by appropriately adjusting the probabilities of ancestral, non-ancestral, or confounding relationships. A detailed discussion is in Appendix F.
5 EXPERIMENTS
Our experiments have three objectives. First, we validate that AGFN approximates well the target distribution over the space of AGs. Second, we show that AGFN performs competitively with alternative methods on both synthetic and real-world datasets. Third, we attest that our experimental design for incorporating the expert’s feedback efficiently reduces the uncertainty over AGFN’s distribution. We provide further experimental details in Appendix C, jointly with results on different datasets, such as DREAM3 Challenge (Marbach et al.,, 2009). Code for reproducing the results were submitted as supplemental material.
5.1 Distributional Assessment of AGFN
Data. As faithfulness violations are more likely in dense graphs (Uhler et al.,, 2012), we generate 2020 55-node random graphs using a directed configuration model (Newman,, 2010), with in- and out-degrees uniformly sampled from $\{0,\ldots,4\}$. For each graph, we draw $500$ samples from a structure-compatible linear Gaussian SCM with randomly chosen parameters.
Setup. We train AGFN for each random graph using their respective samples. Then, we collect AGFN samples to compute empirical distributions over the i) edge features (i.e., $p_{\theta}(V_{i}\rightarrow V_{j})$ and $p_{\theta}(V_{i}\leftrightarrow V_{j})$ for each pair $(V_{i},V_{j})$), ii) path features (i.e., $p_{\theta}(V_{i}\rightsquigarrow V_{j})$) iii) BIC, and iv) Structural Hamming Distance (SHD) to the true causal diagram.
Results. Figure 3 shows that AGFN adequately approximates the theoretical distribution induced by Equation 8. It also induces BIC and SHD distributions that closely match those of $p(\mathcal{G})\propto R(\mathcal{G})$. Here we highlight a critical limitation of N-ADMG, a baseline probabilistic CD method: over $60\%$ of its generated graphs were non-ancestral, severely limiting its utility for inference over AGs. By contrast, AGFN ensures that all generated samples are valid AGs.
5.2 Comparison with SOTA CD algorithms
Data. We generate $10$ datasets with $500$ independent samples from the randomly parametrized linear Gaussian SCMs corresponding to the canonical causal diagram (Richardson and Spirtes,, 2002) of four 4-node AGs with increasingly structural complexity: i) chain4: $W\rightarrow X\rightarrow Y\rightarrow Z$, a simple chain without latent confounders; ii) collfork: $X\rightarrow Z\leftarrow W\rightarrow Y$; with $X\leftrightarrow W$ and $Z\leftrightarrow Y$, forming triplets with colliders and non-colliders under latent confounding, and iii) IV: $X\rightarrow Y$ with $W\rightarrow X\leftarrow Z\rightarrow Y$, a discriminating path for $Z$. Those capture unshielded colliders and discriminating paths, fundamental patterns identifiable by CD algorithms under latent confounding (Spirtes and Richardson,, 1997; Zhang, 2008b, ).
Figure 4: Human-aided AGFN outperforms CD baselines after a single feedback in the considered datasets. Each plot shows the expected SHD under a varying number of feedbacks over $30$ EITL simulations.
Baselines. We compare AGFN with five notable CD methods: FCI (Spirtes et al.,, 2001; Zhang, 2008b, ), GFCI (Ogarrio et al.,, 2016), ACI (Magliacane et al.,, 2016), DCD (Bhattacharya et al.,, 2021), and N-ADMG (Ashman et al.,, 2023), spanning different CD paradigms. FCI is a seminal constraint-based CD algorithm that learns a PAG consistent with conditional independencies entailed by statistical tests. GFCI is a hybrid CD algorithm that learns a PAG by first obtaining an approximate structure using FGS (Ramsey,, 2015) (a BIC-score-based search algorithm for causally sufficient scenarios) and then by applying FCI to identify possible confounding and remove some edges added by FGS. ACI (Magliacane et al.,, 2016) translates conditional independencies into weighted logical statements and uses an Answer Set Programming solver to find an optimal AG. DCD casts CD as continuous optimization with differentiable algebraic constraints defining the space of AGs and uses gradient-based algorithms to solve it. N-ADMG computes a variational approximation of the joint posterior distribution over the space of bow-free causal diagrams (Nowzohour et al.,, 2017) of non-linear SCMs with additive noise. For a more detailed description, we refer to Section C.1.
Experimental setup. We train AGFN on each dataset and sample 100k AGs. For comparison, we apply FCI, GFCI, ACI, and DCD to 100 bootstrapped resamplings of each dataset to approximate the confidence distributions these algorithms induce. Also, we compute the sample mean and standard deviation of BIC and SHD between each method’s estimated PAG and the ground-truth PAG. For methods that output a PAG member (DCD, N-ADMG, and AGFN) we first extract the corresponding PAG before computing the SHD. BIC, is computed directly, as all PAG members are asymptotically score-equivalent.
Results. Summary statistics for the distribution over AGs induced by AGFN and by other baseline CD algorithms are provided in Table 4 in the supplement. Notably, both probabilistic and non-probabilistic methods achieve comparable BIC performance, with AGFN incurring in significantly more diverse samples. Table 1 shows that the three highest-reward samples from AGFN match or exceed the performance of the baselines. For N-ADMG, results comprise the three most frequent samples from its variational distribution.
Table 1: SHD for point estimates. Mean SHD of the AGFN’s top-3 draws is comparable to or better than baselines.